OpenStudy (anonymous):

What is the mean of the following set? 36, 30, 34, 38, 35 a.34.5 b.43.25 c.28.6 d.34.6 PLEase help I do give medals @chosenmatt

OpenStudy (anonymous):

@1497 @micahm @Escobar17

OpenStudy (igreen):

Mean is the average..add all the numbers then divide by how many numbers you have.

OpenStudy (micahm):

The "mean" is the "average" you're used to, where you add up all the numbers and then divide by the number of numbers. The "median" is the "middle" value in the list of numbers.

OpenStudy (anonymous):

I got it all you do is add 36+30+34+38+35= 138 the divde it how many numbers there are, there is 5 so you divide 138/5 and you get 27.6

OpenStudy (igreen):

Actually you added incorrectly.

OpenStudy (anonymous):

36+ 30+ 34+ 38+35=173 173/5=34.6 so D

OpenStudy (igreen):

\(36+30+34+38+35 \neq 138\)

OpenStudy (anonymous):

LOl i hit the wrong button,

OpenStudy (igreen):

@1497 Please don't give answers. We are here to guide the asker to the solution so that they get it on their own.

OpenStudy (anonymous):

whatever u say at least I worked out and not randomly gave the ans